Starting with two coaxial cylinders with a small gap between the inner and outer cylinders and the outer clylinder is fixed while the inner cylinder rotates with an angular velocity w1. The governing equation for the setup is 0 = d/dr ((1/r) (d/dr)(rvthetha)). Show the velocity profile is vthetha(r) = (w1r12)/(r22-r12)*((r22/r )-r)
Answer #1

Given the governing differential equation:

\frac{d}{dr}\left(\frac{1}{r} \frac{d}{dr}(rv_\theta) \right ) = 0\\ \\ \text{Integrating,}\\ \\ \frac{1}{r} \frac{d}{dr}(rv_\theta) = C_1 \\ \\ \text{Integrating again;} \\ \\ v_\theta = \frac{C_1 r}{2} + \frac{C_2}{r}

The boundary conditions are;

\text{At} \,\, r=r_1 \rightarrow v_\theta =\omega r_1\\ \text{At} \,\, r=r_2 \rightarrow v_\theta =0

Now, at r=r_1

\omega r_1 = \frac{C_1 r_1}{2} + \frac{C_2}{r_1} \\ \\ {\color{Red} 2\omega r_1 ^2 = C_1 r_1^2 + 2 C_2}

At, r=r_2

0 = \frac{C_1 r_2}{2} + \frac{C_2}{r_2} \\ \\ {\color{Red} 0= C_1 r_2^2 + 2 C_2}

Solving simultaneously;

{\color{Red} C_1 = \frac{2\omega r_1^2}{(r_1^2 - r_2^2)} \, \text{and}\, \, \\ \\ C_2 = \frac{-\omega r_1^2 r_2^2}{(r_1^2 - r_2^2)}}

On substitution;

v_\theta = \frac{\omega r_1^2 r}{(r_1^2 - r_2^2)} - \frac{\omega r_1^2 r_2^2}{r(r_1^2 - r_2^2)}\\ \\ = \frac{\omega r_1^2 }{(r_1^2 - r_2^2)} \left(r-\frac{r_2^2}{r} \right )

or

v_\theta = \frac{\omega r_1^2 }{(r_2^2 - r_1^2)} \left(\frac{r_2^2}{r} - r\right )